Find the larger of two consecutive integers whose sum is 4 less than three times the smaller integer.

Answer by stanbon(75887)

Find the larger of two consecutive integers whose sum is 4 less than three times the smaller integer.
----------------------
1st: x
2nd: x+1
-----
Equation:
x + x+1 = 3x-4
2x+1 = 3x-4
x = 5
x+1 = 6
==================
Cheers,
Stan h.